OpenStudy (anonymous):

Find two real numbers with a sum of 8 and a product of 12 or show that no such number exists.

Directrix (directrix):

Let x be one of the two real numbers. Because the first number and the second sum to 8, the second number can be written as (8 - x)

Directrix (directrix):

x*(8 - x) = 12 (the product of the two numbers is 12) Solve that equation for x. @jojor_124

OpenStudy (anonymous):

would or could it be something like 6*(8-6) = 12? (8-6 would end up being 2, so 6 times 2 is 12)

Directrix (directrix):

Yes. 6 and 2 multiply to 12 and add to 8.

OpenStudy (anonymous):

Awesome! I think i've got it from here now, thanks!

Directrix (directrix):

You are welcome.